Match the rules to their correct notation: >P(AnB) P(A) (BIA) 1. Addition Rule 2. Multiplication Rule P(A) Independent events Multiplication Rule - Dependent events 3. P(AnB)P(A) P(B) PAUB)PA)P(B)(AnB) 4. Conditional Probability
Answer #1

P(A\cap B)=P(A)*P(B/A)..............3) Multiplication rule for dependent events

P(B/A)=\frac{P(A\cap B)}{P(A)}...................................4).Conditional probability

P(A\cap B)=P(A)*P(B).............................2) Multiplication rule for independent events

P(A\cup B)=P(A)+P(B)-P(A\cap B)......................1)...Addition rule